Question 111 pts

Assume that individuals are homogeneous and that each has a demand curve of the following form for internet service: p=50-2q where p is the price per hour and q is hours per month. Assume the firm has a constant marginal cost of $12. The profit maximising two-part tariff results in the firm setting a per unit price equal to ______ and earning ________ profit from each consumer:

Group of answer choices

12: 361.

12: 589:

31: 361.

31: 589.

None of the above.

Question 121 pts

Consider a monopolist that has two types of consumers. The first, students have a demand curve given by the following: QA=120-2p. The second type of consumer are non-students who have the following demand curve: QB=200-4p. If the monopolist has constant marginal and average total cost equal to 20, which of the following is true if the monopolist practices third degree price discrimination?

Group of answer choices

The price charged to students equals 35 and non-students equals 60.

The price charged to students equals 40 and non-students equals 35.

The price charged to students equals 35 and non-students equals 40.

The price charged to students equals 40 and non-students equals 60.

None of the above

Question 131 pts

Consider a monopolist that has two types of consumers. The first, students have a demand curve given by the following: QA=120-2p. The second type of consumer are non-students who have the following demand curve: QB=200-4p. If the monopolist has constant marginal and average total cost equal to 20, which of the following is true if the monopolist practices third degree price discrimination?

Group of answer choices

The quantity sold to students equals 35 and non-students equals 60.

The quantity sold to students equals 40 and non-students equals 35.

The quantity sold to students equals 35 and non-students equals 40.

The quantity sold to students equals 40 and non-students equals 60.

None of the above

Question 141 pts

Consider the model of monopolistic competition discussed in lectures. If the firm has standard U-shaped cost curves and earns positive economic profit in the short-run which of the following is true in the short run?

Group of answer choices

Price is greater than average total cost.

Price is greater than marginal revenue.

Price is greater than marginal cost.

b and c are correct.

a, b and c are all correct.

Question 151 pts

Consider the model of monopolistic competition discussed in lectures. If the firm has standard U-shaped cost curves and earns zero economic profit in the long-run which of the following is true in the long run?

Group of answer choices

Price is greater than average total cost.

Price is greater than marginal revenue.

Price is greater than marginal cost.

b and c are correct.

a, b and c are all correct.

Question 161 pts

Consider a single price monopolist that faces the following demand curve: p=150-Q. The total cost curve for this monopolist is given by the following: TC=100+10Q+Q2. Which of the following is true?

Group of answer choices

The monopolist will earn profit of -450 and leave the industry in the long run.

The monopolist will earn zero profit and leave the industry in the long run.

The monopolist will earn zero profit but remain in the industry in the long run.

The monopolist will earn profit of 2350 and remain in the industry in the long run.

The monopolist will earn profit of 4025 and remain in the industry in the long run.

Question 171 pts

Bill and Jane are considering buying a mobile phone plan. The retailer offers two types of plans – the standard and the deluxe. Valuations for Bill and Jane for each plan are in the table below:

Bill Jane
Standard 15 25
Deluxe 30 55



Assume that each consumer buys one plan and chooses the plan which maximises their consumer surplus. A plan will be purchased as long as consumer surplus is non-negative. If the marginal cost of each plan is zero, which set of prices for the standard (pS) and deluxe plans (pD ) will maximise profits if the retailer must engage in menu pricing or second degree price discrimination?Group of answer choices

pS=25 and pD=25.

pS=15 and pD=55.

pS=55 and pD=55.

pS=15 and pD=44.

pS=14 and pD=46.

Question 181 pts

Consider the demand curves for beer: Q= 40 –2P. If the supply curve for beer is P=10 and the government imposes a tax of 2 on beer, the total revenue raised is equal to:

Group of answer choices

2

10

12

16

none of the above

Question 191 pts

Assume that demand is perfectly elastic and supply is upward sloping. If a specific tax is imposed on sellers then following the imposition of the tax:

Group of answer choices

Producer surplus is equal to zero.

Consumer surplus is equal to zero.

The total price paid by buyers will increase.

The price received and kept by sellers remains unchanged.

b, c and d are correct.

Question 201 pts

Consider the demand curve for beer: Q= 120 –2P. If the supply curve for beer is P=12+Q and the government imposes a tax of 9 on beer, the deadweight loss that results is equal to:

Group of answer choices

4.5

6

9

27

none of the above
Answer #1

It's mandatory to answer only first Four MCQ

Q11) option A)

12) option B

13) option D

14) option E

15) option D)

M-2, then two part tariff profit = c5 = &[50-19Jx19 p= 50-29, P =MC= 12 12=50-29 29 - 38, q= 19 = (192 = 361 option A) $12,36